# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v1

@@ Appendix B Limitations
Here $\mathbf{k}=(\mathbf{\hat{L}}^{T})^{-1}[\mathbf{b}_{y=1}-\mathbf{b}_{y=0},...,\mathbf{b}_{y=\ell}-\mathbf{b}_{y=0}]$. Further, due to the diversity condition, the matrix $\mathbf{L}=\big{(}p(\mathbf{c}|y=y_{1})-p(\mathbf{c}|y=y_{0}),...,p(\mathbf{c}|y=y_{\ell})-p(\mathbf{c}|y=y_{0})\big{)}$ of size $\ell\times\ell$ is also invertible, $\mathbf{A}$ is invertible.
@@
From the theorem 3.1, we know that:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k}.$ (22)
For two inputs $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$, we have:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x}_{0})}+\mathbf{k}.$ (23)
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x}_{1})}+\mathbf{k}.$ (24)
Subtract the two equations:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})-\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx\mathbf{A}(\log{p(\mathbf{c}|\mathbf{x}_{1})}-\log{p(\mathbf{c}|\mathbf{x}_{0})}).$ (25)
Thus,
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})-\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx\mathbf{A}\triangle\log p(\mathbf{c}),$ (26)
where:
$\triangle\log p(\mathbf{c})=(\log{p(\mathbf{c}|\mathbf{x}_{1})}-\log{p(\mathbf{c}|\mathbf{x}_{0})}).$ (27)
By assumption, $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ differ only in the $i$-th latent variable $c_{i}$, such that:
For $\mathbf{x}_{0}$, $c_{i}=0$, while for $\mathbf{x}_{1}$, $c_{i}=1$. All other $c_{i}$ ($j\neq i$) are the same. This means that the difference in log-probabilities, $\triangle\log p(\mathbf{c})$, will have nonzero contributions only in the $i$-th component. Specifically:
$\triangle\log p(\mathbf{c})=(\log{p(c_{i}=1|\mathbf{x}_{1})}-\log{p(c_{i}=0|\mathbf{x}_{0})}).$ (28)
Define this difference as:
$s=(\log{p(c_{i}=1|\mathbf{x}_{1})}-\log{p(c_{i}=0|\mathbf{x}_{0})}).$ (29)
Thus, $\triangle\log p(\mathbf{c})$ can be written as a vector with only the $i$-th component being nonzero: $\triangle\log p(\mathbf{c})=s\mathbf{e}_{i}$, where $\mathbf{e}_{i}$ is the $i$-th standard basis vector (a vector of zeros with a 1 at the $i$-th position).
Substitute $\triangle\log p(\mathbf{c})$ into Eq. 26, we conclude:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})-\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx s\mathbf{A}_{:,i}.$ (30)
@@
Let $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k}$ as given in Theorem 3.1. For a data pair $(\mathbf{c},\mathbf{x})$, where $\mathbf{x}$ is generated by the latent variable $\mathbf{c}$, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ are passed to a linear classifier with weights $\mathbf{W}$.
@@
$\mathbf{z}\approx\mathbf{W}\big{(}\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k}\big{)}.$ (31)
Since the bias term $\mathbf{k}$ does not affect the classification decision, we focus on the term $\mathbf{W}\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}$. For correct classification, the logits must match the true log-probabilities:
@@
$\mathbf{W}\mathbf{A}=\mathbf{I},$ (33)
which ensures that the classifier produces the correct logits. Here $\mathbf{I}$ denotes the identify matrix. The cross-entropy loss function is minimized when the predicted probabilities match the true probabilities, which is achieved when the logits are correctly aligned with $\log{p(\mathbf{c}|\mathbf{x})}$. Therefore, the classifier weights $\mathbf{W}$ are related to $\mathbf{A}$ by the invertible relation $\mathbf{W}\mathbf{A}=\mathbf{I}$, ensuring that the learned features can be mapped back to the latent distribution, and the classifier can make accurate predictions.
Appendix F Bound of the Jensen Gap
Throughout the proof in Section C, the final approximation of identifiability stems from the application of Jensen’s inequality, specifically Eqs. (11) - (13). Consequently, our focus shifts to analyzing the bound on Jensen’s inequality, commonly referred to as the Jensen gap. Although some prior works have explored this topic in the context of discrete distributions (Simic, 2009, 2008), we recommend that readers interested in the bounds provided by Jensen’s inequality refer to the results from (Gao et al., 2017), as the conditions they impose are weaker and more general (it also can be applied for discrete case). For completeness, we re-write the bound as follows:
Theorem F.1 ((Gao et al., 2017)).
Let $f:\mathcal{I}\to\mathbb{R}$, where $\mathcal{I}$ is a closed subset of $\mathbb{R}$, and let $\mu=\mathbb{E}_{p(\mathbf{c}|y)}[\mathbf{c}]$ be the expectation of $\mathbf{c}$ under the distribution $p(\mathbf{c}|y)$. Suppose the following conditions hold:
1. $f$ is bounded on any compact subset of $\mathcal{I}$: There exists a finite constant $M_{c}$ such that $|f(\mathbf{c})|\leq M_{c}$ for all $\mathbf{c}\in\mathcal{I}$.
2. $|f(\mathbf{c})-f(\mu)|=O(|\mathbf{c}-\mu|^{\alpha})$ as $\mathbf{c}\to\mu$, where $\alpha>0$: This implies that near $\mu$, the deviation of $f(\mathbf{c})$ from $f(\mu)$ grows at most at the rate of $|\mathbf{c}-\mu|^{\alpha}$.
3. $|f(\mathbf{c})|=O(|\mathbf{c}|^{n})$ as $|\mathbf{c}|\to\infty$, where $n\geq\alpha$: This ensures that as $|\mathbf{c}|\to\infty$, $f(\mathbf{c})$ grows at most at the rate of $|\mathbf{c}|^{n}$.
Then, the following inequality holds for the Jensen gap:
$\big{|}\mathbb{E}_{p(\mathbf{c}|y)}[f(\mathbf{c})]-f(\mu)\big{|}\leq M\big{(}\sigma_{\alpha}^{\alpha}+\sigma_{n}^{n}\big{)}\leq M\big{(}1+\sigma_{n}^{n-\alpha}\sigma_{\alpha}^{\alpha}\big{)},$
where:
1. $M$ is a constant defined as:
$M=\sup_{\mathbf{c}\in\mathcal{I}\setminus\{\mu\}}\frac{|f(\mathbf{c})-f(\mu)|}{|\mathbf{c}-\mu|^{\alpha}+|\mathbf{c}-\mu|^{n}}.$
This quantifies the combined growth behavior of $f(\mathbf{c})$ near $\mu$ (captured by $|\mathbf{c}-\mu|^{\alpha}$) and at infinity (captured by $|\mathbf{c}-\mu|^{n}$).
2. $\sigma_{\alpha}^{\alpha}$ is the $\alpha$-th central moment of $\mathbf{c}$ about $\mu$:
$\sigma_{\alpha}^{\alpha}=\mathbb{E}_{p(\mathbf{c}|y)}\big{[}|\mathbf{c}-\mu|^{\alpha}\big{]}.$
3. $\sigma_{n}^{n}$ is the $n$-th raw moment of $\mathbf{c}$:
$\sigma_{n}^{n}=\mathbb{E}_{p(\mathbf{c}|y)}\big{[}|\mathbf{c}|^{n}\big{]}.$
4. $\sigma_{n}^{n-\alpha}\sigma_{\alpha}^{\alpha}$ represents an interaction term between the $n$-th raw moment and the $\alpha$-th central moment:
$\sigma_{n}^{n-\alpha}\sigma_{\alpha}^{\alpha}=\big{(}\mathbb{E}_{p(\mathbf{c}|y)}\big{[}|\mathbf{c}|^{n}\big{]}\big{)}^{(n-\alpha)/n}\big{(}\mathbb{E}_{p(\mathbf{c}|y)}\big{[}|\mathbf{c}-\mu|^{\alpha}\big{]}\big{)}^{\alpha/n}.$
Summary: This inequality bounds the Jensen gap $\big{|}\mathbb{E}_{p(\mathbf{c}|y)}[f(\mathbf{c})]-f(\mu)\big{|}$ using:
  • The growth behavior of $f(\mathbf{c})$ (captured by $M$),
  • The moments of $p(\mathbf{c}|y)$ (captured by $\sigma_{\alpha}^{\alpha}$ and $\sigma_{n}^{n}$).
For more details, refer to (Gao et al., 2017).
Appendix G Experimental Details
G.1 Simulation Details

### v2

@@ Appendix B Limitations
Here due to the diversity condition, the matrix $\mathbf{L}=\big{(}p(\mathbf{c}|y=y_{1})-p(\mathbf{c}|y=y_{0}),...,p(\mathbf{c}|y=y_{\ell})-p(\mathbf{c}|y=y_{0})\big{)}$ of size $\ell\times\ell$ is also invertible, $\mathbf{A}$ is invertible.
Now, we focus on the two terms in the right of Eq. (22), i.e., $\mathbf{{h}}_{y}$ and $\mathbf{{b}}_{y}$, from the viewpoints of when the mapping from latent space, e.g., $\mathbf{c}$, to observed space, e.g., $\mathbf{x}$ and $y$ is invertible and approximately invertible.
When the mapping $\mathbf{g}$ from latent space to observed space is invertible.
We assume that the mapping from the latent space $\mathbf{c}$ to the observed space $(\mathbf{x},y)$ is approximately invertible, meaning that
$1-p(\mathbf{c}=\mathbf{c}^{*}\mid\mathbf{x},y)=\epsilon,$ (23)
where $10\leq\epsilon<$ 1, and $\mathbf{c}^{*}=\arg\max_{\mathbf{c}}p(\mathbf{c}\mid\mathbf{x},y)$ denotes the dominant latent configuration. When the mapping $\mathbf{g}$ is exactly invertible, we have $\epsilon=0$, implying perfect identifiability of the latent variable from the observed data.
  • For the quantity $\mathbf{h}_{y}$, we analyze each component, i.e., $h_{y_{i}}-h_{y_{0}}$, where
    $h_{y_{i}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(\mathbf{c}\mid\mathbf{x},y=y_{i})\right],\quad h_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(\mathbf{c}\mid\mathbf{x},y=y_{0})\right].$ (24)
    When $\epsilon=0$, the posterior distribution $p(\mathbf{c}\mid\mathbf{x},y)$ becomes a delta distribution centered at $\mathbf{c}^{*}$, i.e.,
    $p(\mathbf{c}\mid\mathbf{x},y)=\delta(\mathbf{c}-\mathbf{c}^{*}),$ (25)
    which implies that
    $\log p(\mathbf{c}\mid\mathbf{x},y)=\log\delta(\mathbf{c}-\mathbf{c}^{*}).$ (26)
    In this case, since the posterior concentrates on a single value, its expectation under any distribution $p(\mathbf{c}\mid y)$ becomes constant, and thus
    $h_{y_{i}}-h_{y_{0}}=0.$ (27)
  • Similarly, Recall the definition of the quantity $\mathbf{b}_{y}$. For any pair of labels $y_{i}$ and $y_{0}$, the difference between components is given by:
    $b_{y_{i}}-b_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(y=y_{i}\mid\mathbf{c})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(y=y_{0}\mid\mathbf{c})\right].$ (28)
    Now since the mapping from latent variable $\mathbf{c}$ to observed pair $(\mathbf{x},y)$ is invertible, whch implies that for each $y=y_{k}$, the conditional distribution $p(\mathbf{c}\mid y=y_{k})$ is supported only on values of $\mathbf{c}$ such that $p(y=y_{k}\mid\mathbf{c})=1$. That is:
    $p(y=y_{k}\mid\mathbf{c})=1\quad\text{for }\mathbf{c}\sim p(\mathbf{c}\mid y=y_{k}).$ (29)
    Therefore, for any such $\mathbf{c}$, we have:
    $\log p(y=y_{k}\mid\mathbf{c})=\log 1=0.$ (30)
    Hence, each term in the difference becomes:
    $\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\log p(y=y_{i}\mid\mathbf{c})=0,$ (31)
    $\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\log p(y=y_{0}\mid\mathbf{c})=0.$ (32)
    Thus, the difference is:
    $b_{y_{i}}-b_{y_{0}}=0-0=0.$ (33)
When the mapping $\mathbf{g}$ is approximately invertible, i.e., $\epsilon\rightarrow 0$,
  • For the quantity $\mathbf{h}_{y}$, We analyze the component difference:
    $h_{y_{i}}-h_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(\mathbf{c}\mid\mathbf{x},y_{i})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(\mathbf{c}\mid\mathbf{x},y_{0})\right].$
    Since the generative map $\mathbf{g}:\mathbf{c}\mapsto(\mathbf{x},y)$ is approximately invertible, the posterior $p(\mathbf{c}\mid\mathbf{x},y)$ becomes sharply concentrated at a unique $\mathbf{c}^{*}(\mathbf{x},y)$. That is, for any small $\epsilon>0$, the posterior takes the form:
    $p(\mathbf{c}\mid\mathbf{x},y)=(1-\epsilon)\cdot\delta_{\mathbf{c}^{*}}+\epsilon\cdot q(\mathbf{c}),$
    where $\delta_{\mathbf{c}^{*}}$ is a point mass at $\mathbf{c}^{*}(\mathbf{x},y)$, and $q(\mathbf{c})$ is a distribution over other latent codes with bounded support and $\max_{\mathbf{c}\neq\mathbf{c}^{*}}\log p(\mathbf{c}\mid\mathbf{x},y)\leq\log\epsilon$.
    Then for each expectation:
    $\mathbb{E}_{p(\mathbf{c}\mid y)}\left[\log p(\mathbf{c}\mid\mathbf{x},y)\right]=p(\mathbf{c}^{*}\mid y)\log(1-\epsilon)+\sum_{\mathbf{c}\neq\mathbf{c}^{*}}p(\mathbf{c}\mid y)\log p(\mathbf{c}\mid\mathbf{x},y).$
    Now note:
    • $\log(1-\epsilon)\approx 0$,
    • $\log p(\mathbf{c}\mid\mathbf{x},y)\leq\log\epsilon$ for all $\mathbf{c}\neq\mathbf{c}^{*}$,
    • and for $\mathbf{c}\neq\mathbf{c}^{*}$ $p(\mathbf{c}\mid y)$ is bounded.
    So the entire expectation satisfies:
    $\mathbb{E}_{p(\mathbf{c}\mid y)}\left[\log p(\mathbf{c}\mid\mathbf{x},y)\right]\approx\log\epsilon,$
    which do not depend on $y$. Therefore,
    $h_{y_{i}}-h_{y_{0}}\xrightarrow[\epsilon\to 0]{}0.$
  • Recall the definition of the quantity $\mathbf{b}_{y}$. For any pair of labels $y_{i}$ and $y_{0}$, the difference between components is:
    $b_{y_{i}}-b_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(y=y_{i}\mid\mathbf{c})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(y=y_{0}\mid\mathbf{c})\right].$ (34)
    Since the generative process $\mathbf{g}:\mathbf{c}\rightarrow(\mathbf{x},y)$ is approximately invertible, i.e., $\epsilon\to 0$, the conditional $p(y\mid\mathbf{c})$ becomes sharply peaked. In particular, there exists a dominant latent mode $\mathbf{c}^{*}_{y_{i}}$ such that
    $p(y=y_{i}\mid\mathbf{c}^{*}_{y_{i}})=1-\epsilon^{\prime},\quad p(y\neq y_{i}\mid\mathbf{c}^{*}_{y_{i}})=\epsilon^{\prime},$
    where $\epsilon^{\prime}\to 0$ as $\epsilon\to 0$.
    Therefore, the expected log-probability becomes:
    $\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(y=y_{i}\mid\mathbf{c})\right]$ $=p(\mathbf{c}^{*}_{y_{i}}\mid y=y_{i})\log p(y=y_{i}\mid\mathbf{c}^{*}_{y_{i}})$
    $\quad+\sum_{\mathbf{c}^{\prime}\neq\mathbf{c}^{*}_{y_{i}}}p(\mathbf{c}^{\prime}\mid y=y_{i})\log p(y=y_{i}\mid\mathbf{c}^{\prime})$ (35)
    $\approx\log(1-\epsilon^{\prime})+\sum_{\mathbf{c}^{\prime}\neq\mathbf{c}^{*}_{y_{i}}}p(\mathbf{c}^{\prime}\mid y=y_{i})\log\epsilon^{\prime}.$ (36)
    As $\epsilon^{\prime}\to 0$, we have $\log(1-\epsilon^{\prime})\to 0$. Further, since $p(\mathbf{c}^{\prime}\mid y=y_{i})$ is bounded, the second tends to be $\log\epsilon^{\prime}$, which only depends on $\epsilon^{\prime}$. Therefore, the difference becomes:
    $b_{y_{i}}-b_{y_{0}}\xrightarrow[\epsilon\to 0]{}0.$
@@
According to the assumption in Theorem 3.1, there exists a matrix $\mathbf{A}$ such that:
$\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}$ (37)
where $[\cdot]_{\mathbf{c}_{i}}$ denotes concatenating $\log p(\mathbf{c}_{i}|\mathbf{x})$ over all possible values of $\mathbf{c}$.
Since $1-p(\mathbf{c}=\mathbf{c}^{*}_{\mathbf{x}}|\mathbf{x})=\epsilon_{\mathbf{x}}$, when $\epsilon_{\mathbf{x}}\rightarrow 0$, we have:
$\left[\log p(\mathbf{c}={\mathbf{c}_{i}}|\mathbf{x})\right]_{\mathbf{c}_{i}}\approx[\log o(\epsilon_{\mathbf{x}}),...,\log(1-\epsilon_{\mathbf{x}}),...,\log o(\epsilon_{\mathbf{x}})],$ (38)
where where $\log(1-\epsilon_{x})$ appears at the index corresponding to $\mathbf{c}_{x}^{*}$, and other entries are $\log o(\epsilon_{x})$.
Now consider two inputs $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$, with their most likely concepts being $\mathbf{c}_{x_{0}}^{*}$ and $\mathbf{c}_{x_{1}}^{*}$, corresponding to indices $k_{0}$ and $k_{1}$, respectively, with $k_{0}\neq k_{1}$. We have:
$f_{\mathbf{x}}(\mathbf{x}_{1})-f_{\mathbf{x}}(\mathbf{x}_{0})$ $\approx\mathbf{A}\left(\left[\log p(\mathbf{c}=\mathbf{c}_{j}\mid\mathbf{x}_{1})\right]_{\mathbf{c}_{j}}-\left[\log p(\mathbf{c}=\mathbf{c}_{j}\mid\mathbf{x}_{0})\right]_{\mathbf{c}_{j}}\right)$ (39)
$\approx\mathbf{A}\left(\left[0,\ldots,\underbrace{\log(1-\epsilon_{\mathbf{x}_{1}})-\log o(\epsilon_{\mathbf{x}_{0}})}_{k_{1}},\ldots,\underbrace{\log o(\epsilon_{\mathbf{x}_{1}})-\log(1-\epsilon_{\mathbf{x}_{0}})}_{k_{0}},\ldots,0\right]\right)$ (40)
Only the $k_{1}$-th and $k_{0}$-th positions are nonzero, so this can be written as:
$f_{\mathbf{x}}(\mathbf{x}_{1})-f_{\mathbf{x}}(\mathbf{x}_{0})\approx\left(\underbrace{\log(1-\epsilon_{\mathbf{x}_{1}})-\log o(\epsilon_{\mathbf{x}_{0}})}_{s_{1}}\right)\mathbf{A}_{:,k_{1}}+\left(\underbrace{\log o(\epsilon_{\mathbf{x}_{1}})-\log(1-\epsilon_{\mathbf{x}_{0}})}_{s_{2}}\right)\mathbf{A}_{:,k_{0}}$
Then we obtain:
$f_{\mathbf{x}}(\mathbf{x}_{1})-f_{\mathbf{x}}(\mathbf{x}_{0})\approx s_{1}\mathbf{A}_{:,k_{1}}-s_{2}\mathbf{A}_{:,k_{0}}$
Since the inputs $\mathbf{x}_{1}$ and $\mathbf{x}_{0}$ differ only in the $i$-th concept variable $c_{i}$, and share identical values for all other concept variables $\mathbf{c}_{-i}$. Specifically, let:
  • $\mathbf{x}_{1}$ corresponds to concept assignment $\mathbf{c}_{1}^{*}=(c_{i}=1,\mathbf{c}_{-i}^{*})$;
  • $\mathbf{x}_{0}$ corresponds to concept assignment $\mathbf{c}_{0}^{*}=(c_{i}=0,\mathbf{c}_{-i}^{*})$;
  • The conditional probabilities satisfy $p(\mathbf{c}_{1}^{*}\mid\mathbf{x}_{1})=1-\epsilon_{\mathbf{x}_{1}}$ and $p(\mathbf{c}_{0}^{*}\mid\mathbf{x}_{0})=1-\epsilon_{\mathbf{x}_{0}}$, where $\epsilon_{\mathbf{x}_{1}},\epsilon_{\mathbf{x}_{0}}\to 0$.
Now consider the conditional probability $p(c_{i}=1\mid\mathbf{x}_{1})$:
$p(c_{i}=1\mid\mathbf{x}_{1})=\sum_{\mathbf{c}_{-i}}p(c_{i}=1,\mathbf{c}_{-i}\mid\mathbf{x}_{1})=p(\mathbf{c}_{1}^{*}\mid\mathbf{x}_{1})+\sum_{\mathbf{c}_{-i}^{\prime}\neq\mathbf{c}_{-i}^{*}}p(c_{i}=1,\mathbf{c}_{-i}^{\prime}\mid\mathbf{x}_{1}).$
Because $p(\mathbf{c}_{1}^{*}\mid\mathbf{x}_{1})=1-\epsilon_{\mathbf{x}_{1}}$ and the remaining probability mass is small (order $\epsilon_{\mathbf{x}_{1}}$), we obtain:
$p(c_{i}=1\mid\mathbf{x}_{1})\approx 1-\epsilon_{\mathbf{x}_{1}}.$
Thus:
$s_{1}=\log(1-\epsilon_{\mathbf{x}_{1}})-\log o(\epsilon_{\mathbf{x}_{0}})\approx\log(1-\epsilon_{\mathbf{x}_{1}})\approx\log p(c_{i}=1\mid\mathbf{x}_{1}),$
and similarly,
$s_{2}\approx\log p(c_{i}=0\mid\mathbf{x}_{0}).$
@@
Let $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$ as given in Theorem 3.1. For a data pair $(\mathbf{c},\mathbf{x})$, where $\mathbf{x}$ is generated by the latent variable $\mathbf{c}$, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ are passed to a linear classifier with weights $\mathbf{W}$.
@@
$\mathbf{z}\approx\mathbf{W}\big{(}\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}\big{)}.$ (41)
For correct classification under cross entropy loss, the logits must match the true log-probabilities:
@@
$\mathbf{W}\mathbf{A}\approx\mathbf{I},$ (43)
which ensures that the classifier produces the correct logits. Here $\mathbf{I}$ denotes the identify matrix. The cross-entropy loss function is minimized when the predicted probabilities match the true probabilities, which is achieved when the logits are correctly aligned with $\log{p(\mathbf{c}|\mathbf{x})}$. Therefore, the classifier weights $\mathbf{W}$ are related to $\mathbf{A}$ by the invertible relation $\mathbf{W}\mathbf{A}\approx\mathbf{I}$, ensuring that the learned features can be mapped back to the latent distribution, and the classifier can make accurate predictions.
Appendix F Experimental Details
F.1 Simulation Details